OpenStudy (anonymous):

A gift basket has 2 pounds of cookies, 3/5 pound of fudge, and 1 1/4 pound of nuts. what is the total weight of the items in the basket?

OpenStudy (anonymous):

3 5/20= 3 1/4LBS

OpenStudy (anonymous):

2 + 3/5 + 1 1/4 = 2/1 + 3/5 + 5/4 = 40/20 + 12/20 + 25/20 = 77/20 = 3 17/20 lbs